# When Do Credal Sets Stabilize? Fixed-Point Theorems for Credal Set Updates

### v1

@@ Abstract.
Many machine learning algorithms rely on iterative updates of uncertainty representations, ranging from variational inference and expectation-maximization, to reinforcement learning, continual learning, and multi-agent learning. In the presence of imprecision and ambiguity, credal sets—closed, convex sets of probability distributions—have emerged as a popular framework for representing imprecise probabilistic beliefs. Under such imprecision, many learning problems in imprecise probabilistic machine learning (IPML) may be viewed as processes involving successive applications of update rules on credal sets. This naturally raises the question of whether this iterative process converges to stable fixed points—or, more generally, under what conditions on the updating mechanism such fixed points exist, and whether they can be attained. We provide the first analysis of this problem and illustrate our findings using Credal Bayesian Deep Learning as a concrete example. Our work demonstrates that incorporating imprecision into the learning process not only enriches the representation of uncertainty, but also reveals structural conditions under which stability emerges, thereby offering new insights into the dynamics of iterative learning under imprecision.
@@ 1. Introduction
In learning problems, a fixed-point refers to a solution that remains unchanged when a specific transformation or algorithm is applied repeatedly. A fixed-point’s existence often acts as a guarantee of stability for an iterative process, meaning the system’s dynamics will eventually converge to a stable solution (Banach,, 1922; Ortega and Rheinboldt,, 2000). By the Banach Fixed Point Theorem (Kamalov and Leung,, 2023, Theorem 2.1), it is well-known that a continuous function on a complete metric space has a unique fixed point if it is a contraction (see Sections 2 and 3 for technical details). Fixed-point theorems have been used in proving that a learning or optimisation problem has a solution, e.g., an equilibrium point, stable model, or policy (Banach,, 1922; Kakutani,, 1941), showing that an iterative learning rule, e.g., policy update, message passing, and Expectation Maximisation (EM) step, converges to a stable point (Dempster et al.,, 1977; Bertsekas and Tsitsiklis,, 1996), and characterising Nash equilibria or best-response dynamics of agents’ strategies in multi-agent learning (Nash,, 1950; Fudenberg and Tirole,, 1991). Recently, fixed-point analysis has become a valuable tool for understanding the dynamics that arise from the interaction between human and AI systems. For example, it helps researchers analyse phenomena like model collapse (Shumailov et al.,, 2024) and performativity (Perdomo et al.,, 2020; Hardt and Mendler-Dünner,, 2025) that emerge from these complex interactions.
@@
Classical fixed-point analysis, though widely applied, often falls short of capturing the intrinsic ambiguity in learning problems. [^1] This work focuses on the fixed-point analysis of update rules applied on a space of probability distributions. In the domain of continual learning, where models are trained iteratively on a stream of data, the learning dynamics frequently exhibit instability (Kirkpatrick et al.,, 2017; Parisi et al.,, 2019). Empirical observations reveal that the training process often fails to converge to a single, stable solution. This lack of convergence manifests as significant fluctuations in the model’s parameters, which can substantially degrade performance on previously learned tasks. This phenomenon, widely known as catastrophic forgetting, represents a key challenge in developing robust continual learning systems. Furthermore, a fixed-point for iterative update rules may not be attainable when the underlying data sources contain inherent conflicts (Arrow,, 1950; Walley,, 1991; Genest and Zidek,, 1986). These conflicts, which can arise from disparate sources like heterogeneous datasets or conflicting individual opinions, introduce a dynamic instability that prevents convergence to a stable solution.
@@
Inspired by these challenges, this paper develops fixed-point theorems for update rules on credal sets—closed, convex sets of probability distributions (Levi,, 1980; Walley,, 2000)—which remain comparatively understudied relative to fixed-point analysis on the space of single probability distributions. Update rules on credal sets play a central role in imprecise probabilistic machine learning (IPML), a growing field that leverages imprecise probability (IP) theory to enhance robustness, trustworthiness, and safety in machine learning. Classical examples include the generalised Bayes rule (Walley,, 1991), Dempster’s rule (Dempster,, 1967; Shafer,, 1976), the geometric rule (Suppes and Zanotti,, 1977), and others (Caprio and Seidenfeld,, 2023), all of which operate directly on credal sets. More recently, modern approaches such as Credal Bayesian Deep Learning (CBDL) (Caprio et al., 2024a, ) and Imprecise Bayesian Continual Learning (IBCL) (Lu et al.,, 2024) have applied these rules in the context of Bayesian deep learning (BDL) and continual learning (CL). Their historical significance, combined with their increasing adoption in contemporary machine learning, underscores the importance of establishing fundamental conditions under which the stable solutions exist.
@@ 2.1. Fixed Point Theorems
Fixed point theory (Granas et al.,, 2003) studies the existence, uniqueness, and qualitative properties of solutions to equations of the form $f(x)=x$, for a given mapping $f$. Its applications are widespread across disciplines. In economics, fixed point results underpin the existence of equilibria in game theory and general equilibrium models (Scarf,, 1983). In physics, they are used to analyse the long-term behaviour of dynamical systems and phase transitions (Hess and Dancer,, 1991). In computer science and logic, fixed points provide the foundation for induction principles and recursive definitions, ensuring that self-referential constructions are mathematically well grounded (Abel,, 2012).
In probability theory and statistics, fixed point theorems are crucial for establishing the existence of invariant probability measures. Examples include stationary distributions of Markov chains and equilibrium laws in mean-field models. Here, the space of probability measures—convex, compact under weak topologies, and often metrised by distances such as Wasserstein—provides a natural setting in which classical results such as the Banach’s (Banach,, 1922), Schauder’s (Schauder,, 1930), and Kakutani’s (Kakutani,, 1941) fixed point theorems can be applied to prove both existence and uniqueness of such invariant distributions.
However, these classical results are typically formulated for mappings acting on precise probability measures or on compact convex sets of probability vectors in $\mathbb{R}^{d}$. Their applicability becomes less clear when considering mappings on the space of credal sets (Levi,, 1980)—that is, nonempty, weak-closed, convex subsets of probability measures. Such sets extend beyond the finite-dimensional setting of probability vectors and introduce additional structural complexity. Investigating how fixed point results can be generalised to this richer setting is the focus of Section 3. Pursuing this direction promises to provide a rigorous theoretical foundation for imprecise probabilistic machine learning, introduced in the following subsection, where credal sets play a central role.
@@ 2.2. Imprecise Probabilistic Machine Learning
Imprecise Probabilistic Machine Learning (IPML) is an emerging field that seeks to integrate the rigorous theory of imprecise probabilities (Walley,, 1991) into probabilistic machine learning. Its ultimate goal is to be precise about imprecision, thereby enabling more robust and trustworthy models and inferences. Among the various representations of imprecision, the credal set—a nonempty, closed, and convex set of probability measures—is arguably the most central object of study. In recent years, the incorporation of credal sets into machine learning has led to significant developments across a wide range of areas, including classification (Caprio et al., 2024b, ), hypothesis testing (Chau et al., 2025b, ; Jürgens et al.,, 2025), scoring rules (Fröhlich and Williamson,, 2024; Singh et al.,, 2025), conformal prediction (Caprio et al.,, 2025; Caprio,, 2025), computer vision (Cuzzolin and Frezza,, 1999; Giunchiglia et al.,, 2023), probabilistic programming (Liell-Cock and Staton,, 2025), explainability (Chau et al.,, 2023; Utkin et al.,, 2025), neural networks (Caprio et al., 2024a, ; Wang et al.,, 2024, 2025), learning theory (Caprio et al., 2024c, ), causal inference (Cozman,, 2000; Zaffalon et al.,, 2023), active and continual learning (Dutta et al.,, 2024; Lu et al.,, 2024), and probability metrics (Chau et al., 2025a, ), among others.
However, to the best of our knowledge, there has been no prior work studying fixed point theorems—a fundamental and widely applied concept in mathematics—for mappings acting on credal sets. Filling this gap is the main focus of our contribution.
@@ 3. Main Results
Let $\mathcal{X}$ be a measurable space of interest, and call $\Delta_{\mathcal{X}}$ the space of countably additive probability measures on $\mathcal{X}$. While usually in Imprecise Probability theory scholars work with finitely additive probabilities, in this section we focus on the countably additive case to avoid complications. [^2] Notice that, if $\mathcal{X}$ is finite, the results that we find hold also in the finitely additive case (and we also have that $\mathcal{X}$ is always compact as a bonus). Let $\mathscr{C}\subset 2^{\Delta_{\mathcal{X}}}$ be the space of nonempty weak-closed and convex subsets of $\Delta_{\mathcal{X}}$. Let $\mathcal{P}$ be a generic credal set on $\mathcal{X}$ (Levi,, 1980), so that $\mathcal{P}\in\mathscr{C}$.
Remark 1 (On Finite Additivity).
The reason why the results in this section may not hold in the finitely additive case is that, if $\mathcal{X}$ is not finite (see Appendix B), then $\Delta_{\mathcal{X}}$ may not be metrizable, and so we cannot talk about Hausdorff distance (between the elements of $\mathscr{C}$), which plays a crucial role in our proofs.
@@
Let $\mathcal{X}=(\mathcal{X},d)$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is continuous in the Hausdorff topology, [^3] Recall that $f$ is Hausdorff continuous if, for a net $(K_{\gamma})$ in $\mathscr{C}$, $K_{\gamma}\xrightarrow[]{d_{H}}K\implies f(K_{\gamma})\xrightarrow[]{d_{H}}f(K)$, where $d_{H}$ denotes the Hausdorff metric. then the set of fixed points of $f$ is compact and nonempty.
Theorem 1 tells us that, if our space $\mathcal{X}$ of interest has a sufficiently well-behaved topological structure, then a continuous updating function is guaranteed to have (at least) a fixed point. It can be seen as a generalization of Kakutani’s fixed point result (Kakutani,, 1941), declined for credal sets.
Notice that the assumption in Theorem 1 that $\mathcal{X}$ is compact is not overly strong: For example, the space of colored images $0,1]^{w\times h\times 3}$ (where $w$ denotes the images’ width, $h$ their height, and $3$ stands for the RGB option) is compact. In practice, if $\mathcal{X}$ is not compact (in the topology compatible with the metric $d$ that we choose), we can always either change the metric (and hence also its compatible topology), or consider the compactification of $\mathcal{X}$. The real heavy lifting is done by the updating function $f$ being continuous (and assumed to output credal sets). This is a property at the very heart of the updating rule, and it cannot be arbitrarily modified. We now provide two examples, one showing that without continuity, the existence of fixed points is indeed in jeopardy, and one that shows that a popular Imprecise Probabilistic Machine Learning method can be written as a continuous updating mechanism.
@@
  • Discontinuity (Hausdorff). Fix any $a\in[0,1-\delta]$ and let $I_{n}=[a,1-\delta-1/n]\to I=[a,1-\delta]$ in $d_{H}$. Then $f(I_{n})=[a+\delta,1-1/n]\to[a+\delta,1]$, while $f(I)=[\max\{0,a-\delta\},1-\delta]$. Hence $d_{H}\big(f(I_{n}),f(I)\big)\geq\delta$, so $f$ is not Hausdorff-continuous at $\lambda_{2}=1-\delta$.
  • No fixed points. If $\lambda_{2}<1-\delta$, then $f$ shifts the interval right by $\delta$, so equality would force $\delta=0$, impossible. If $\lambda_{2}\geq 1-\delta$, then $f([\lambda_{1},\lambda_{2}])=[\max\{0,\lambda_{1}-\delta\},\lambda_{2}-\delta]$, so equality would require $\lambda_{2}=\lambda_{2}-\delta$, again impossible. Thus, the set $\mathrm{Fix}(f)$ of fixed points of $f$ is empty.
@@
Example 2 (Credal Bayesian Deep Learning, CBDL).

### v2

@@ Abstract.
Many machine learning algorithms rely on iterative updates of uncertainty representations, ranging from variational inference and expectation-maximization, to reinforcement learning, continual learning, and multi-agent learning. In the presence of imprecision and ambiguity, credal sets—closed, convex sets of probability distributions—have emerged as a popular framework for representing imprecise probabilistic beliefs. Under such imprecision, many learning problems in imprecise probabilistic machine learning (IPML) may be viewed as processes involving successive applications of update rules on credal sets. This naturally raises the question of whether this iterative process converges to stable fixed points—or, more generally, under what conditions on the updating mechanism such fixed points exist, and whether they can be attained. We provide the first analysis of this problem, and illustrate our findings using Credal Bayesian Deep Learning as a concrete example. Our work demonstrates that incorporating imprecision into the learning process not only enriches the representation of uncertainty, but also reveals structural conditions under which stability emerges, thereby offering new insights into the dynamics of iterative learning under imprecision.
@@ 1. Introduction
In learning problems, a fixed point refers to a solution that remains unchanged when a specific transformation or algorithm is applied repeatedly. A fixed point’s existence often acts as a guarantee of stability for an iterative process, meaning the system’s dynamics will eventually converge to a stable solution (Banach,, 1922; Ortega and Rheinboldt,, 2000). Fixed point theorems have been used in proving that a learning or optimization problem has a solution—e.g. an equilibrium point, stable model, or policy (Banach,, 1922; Kakutani,, 1941)—showing that an iterative learning rule—e.g. policy update, message passing, and Expectation Maximization (EM) step—converges to a stable point (Dempster et al.,, 1977; Bertsekas and Tsitsiklis,, 1996), and characterizing Nash equilibria or best-response dynamics of agents’ strategies in multi-agent learning (Nash,, 1950; Fudenberg and Tirole,, 1991). Recently, fixed point analysis has become a valuable tool for understanding the dynamics that arise from the interaction between human and AI systems—such as phenomena like model collapse (Shumailov et al.,, 2024) and performativity (Perdomo et al.,, 2020; Hardt and Mendler-Dünner,, 2025) that emerge from these complex interactions—as well as reciprocal learning where multiple agents, models, or systems adapt to each other’s behavior in a mutually dependent way (Rodemann et al.,, 2024; Rodemann and Bailie,, 2025).
@@
Classical fixed point analysis, though widely applied, often falls short of capturing the intrinsic ambiguity in learning problems. In the domain of continual learning, where models are trained iteratively on a stream of data, the learning dynamics frequently exhibit instability (Kirkpatrick et al.,, 2017; Parisi et al.,, 2019). Empirical observations reveal that the training process often fails to converge to a single, stable solution. This lack of convergence manifests as significant fluctuations in the model’s parameters, which can substantially degrade performance on previously learned tasks. This phenomenon, widely known as catastrophic forgetting, represents a key challenge in developing robust continual learning systems. Furthermore, a fixed point for iterative update rules may not be attainable when the underlying data sources contain inherent conflicts (Arrow,, 1950; Walley,, 1991; Genest and Zidek,, 1986). These conflicts, which can arise from disparate sources like heterogeneous datasets or conflicting individual opinions, introduce a dynamic instability that prevents convergence to a stable solution.
@@
Inspired by these challenges, this paper develops fixed point theorems for update rules on credal sets—closed, convex sets of probability distributions (Levi,, 1980; Walley,, 2000)—which remain comparatively understudied relative to fixed point analysis on the space of single probability distributions. Update rules on credal sets play a central role in imprecise probabilistic machine learning (IPML), a growing field that leverages imprecise probability (IP) theory to enhance robustness, trustworthiness, and safety in machine learning. Classical examples include the generalized Bayes rule (Walley,, 1991), Dempster’s rule (Dempster,, 1967; Shafer,, 1976), and the geometric rule (Suppes and Zanotti,, 1977), all of which operate directly on credal sets. More recently, modern approaches such as Credal Bayesian Deep Learning (CBDL) (Caprio et al., 2024a, ) and Imprecise Bayesian Continual Learning (IBCL) (Lu et al.,, 2025) have applied these rules in the context of Bayesian deep learning (BDL) and continual learning (CL). Their historical significance, combined with their increasing adoption in contemporary machine learning, underscores the importance of establishing fundamental conditions under which the stable solutions exist.
@@
Given the page limitations, we provide the necessary mathematical background—namely the definitions of compact metric space, weak topology (in this paper, we call weak-convergence the convergence in expectation of all bounded continuous functions), dominating measure, and likelihood kernel, as well as the statements of Kakutani’s, Edelstein’s, and Boyd-Wong’s fixed point theorems—in Appendix A, and the proofs of our results in Appendix B.
@@ 2.1. Fixed Point Theorems
Fixed point theory (Granas et al.,, 2003) studies the existence, uniqueness, and qualitative properties of solutions to equations of the form $f(x)=x$, for a given mapping $f$. Its applications are widespread across disciplines. In economics, fixed point results underpin the existence of equilibria in game theory and general equilibrium models (Scarf,, 1983). In physics, they are used to analyze the long-term behavior of dynamical systems and phase transitions (Dancer and Hess,, 1991). In computer science and logic, fixed points provide the foundation for induction principles and recursive definitions, ensuring that self-referential constructions are mathematically well grounded (Abel,, 2012).
In probability theory and statistics, fixed point theorems are crucial for establishing the existence of invariant probability measures. Examples include stationary distributions of Markov chains and equilibrium laws in mean-field models. Here, the space of probability measures provides a natural setting in which classical results such as the Banach’s (Banach,, 1922), Schauder’s (Schauder,, 1930), and Kakutani’s (Kakutani,, 1941) fixed point theorems can be applied to prove both existence and uniqueness of such invariant distributions.
However, when viewed from a probabilistic perspective, these classical results are typically formulated for mappings acting on precise probability measures. Their applicability becomes less clear when considering mappings on the space of credal sets (Levi,, 1980)—that is, nonempty, weak-closed, convex subsets of probability measures. Such sets extend beyond the finite dimensional setting of probability vectors and introduce additional structural complexity. Investigating how fixed point results can be generalized to this richer setting is the focus of Section 3. Pursuing this direction promises to provide a rigorous theoretical foundation for imprecise probabilistic machine learning, introduced in the following subsection, where credal sets play a central role.
It is worth noting that the field of imprecise stochastic processes—especially the work on the convergence of imprecise Markov chains (de Cooman et al.,, 2008, 2009; Škulj and Hable,, 2009)—addresses questions that are similar in spirit to those studied in this paper, although typically restricted to (at most) countable state spaces. In particular, their results rely on the Perron-Frobenius theorem, which we conjecture could also be derived using our fixed point methods.
@@ 2.2. Imprecise Probabilistic Machine Learning
Imprecise Probabilistic Machine Learning (IPML) is an emerging field that seeks to integrate the rigorous theory of imprecise probabilities (Walley,, 1991) into probabilistic machine learning. Its ultimate goal is to be precise about imprecision, thereby enabling more robust and trustworthy models and inferences. Among the various representations of imprecision, the credal set is arguably the most central object of study. In recent years, the incorporation of credal sets into machine learning has led to significant developments across a wide range of areas, including classification (Caprio et al., 2024b, ), domain generalization (Singh et al.,, 2024), hypothesis testing (Chau et al., 2025b, ; Jürgens et al.,, 2025), scoring rules (Fröhlich and Williamson,, 2024; Singh et al.,, 2025), conformal prediction (Caprio et al.,, 2025; Caprio,, 2025), computer vision (Cuzzolin and Frezza,, 1999; Giunchiglia et al.,, 2023), probabilistic programming (Liell-Cock and Staton,, 2025), explainability (Chau et al.,, 2023; Utkin et al.,, 2025; Mohammadi et al.,, 2026), neural networks (Caprio et al., 2024a, ; Wang et al.,, 2024, 2025), learning theory (Caprio et al., 2024c, ), causal inference (Cozman,, 2000; Zaffalon et al.,, 2023), active and continual learning (Dutta et al.,, 2024; Lu et al.,, 2025), and probability metrics (Chau et al., 2025a, ), among others.
However, to the best of our knowledge, there has been no prior work studying fixed point theorems for mappings acting on credal sets. Filling this gap is the main focus of our contribution.
@@ 3. Main Results
Let $\mathcal{X}$ be a measurable space of interest, and call $\Delta_{\mathcal{X}}$ the space of countably additive probability measures on $\mathcal{X}$. While usually in Imprecise Probability theory scholars work with finitely additive probabilities, in this section we focus on the countably additive case to avoid complications. The finitely additive case is studied in Appendix D. Let $\mathscr{C}\subset 2^{\Delta_{\mathcal{X}}}$ be the space of nonempty weak-closed and convex subsets of $\Delta_{\mathcal{X}}$. Let $\mathcal{P}$ be a generic credal set on $\mathcal{X}$ (Levi,, 1980), so that $\mathcal{P}\in\mathscr{C}$.
3.1. Existence of a Fixed Point
@@
Recall that a function $f:\mathscr{C}\rightarrow\mathscr{C}$ is Hausdorff continuous if, for a net $(\mathcal{P}_{\gamma})$ in $\mathscr{C}$, $\mathcal{P}_{\gamma}\xrightarrow[]{d_{H}}\mathcal{P}$ implies $f(\mathcal{P}_{\gamma})\xrightarrow[]{d_{H}}f(\mathcal{P})$, where $d_{H}$ denotes the Hausdorff metric. The latter is defined as $d_{H}(\mathcal{P},\mathcal{Q})\coloneqq\max\{\sup_{P\in\mathcal{P}}\rho(P,\mathcal{Q}),\sup_{Q\in\mathcal{Q}}\rho(\mathcal{P},Q)\}$, where $\rho(P,\mathcal{Q})\coloneqq\inf_{Q\in\mathcal{Q}}\rho(P,Q)$ and similarly for $\rho(\mathcal{P},Q)$, and $\rho$ is any metric that metrizes the weak topology, like the Prokhorov metric (Billingsley,, 2013, Theorem 6.8).
@@
Let $\mathcal{X}=(\mathcal{X},d)$ be a compact metric space, and endow $\Delta_{\mathcal{X}}$ with the weak topology. Consider a function $f:\mathscr{C}\rightarrow\mathscr{C}$ that represents our updating technique. If $f$ is continuous in the Hausdorff topology, then the set of fixed points of $f$ is compact and nonempty.
Theorem 1, proved in Appendix B.1, tells us that, if our space $\mathcal{X}$ of interest has a sufficiently well-behaved topological structure, then a continuous updating function is guaranteed to have (at least) a fixed point. Theorem 1 bears similarities to a version of Kakutani’s Fixed Point Theorem, presented in Appendix A.5.
Notice that the assumption in Theorem 1 that $\mathcal{X}$ is compact is not overly strong: For example, the space of colored images $0,1]^{w\times h\times 3}$ (where $w$ denotes the images’ width, $h$ their height, and $3$ stands for the RGB option) is compact. In practice, if $\mathcal{X}$ is not compact (in the topology compatible with the metric $d$ that we choose), we can always either change the metric (and hence also its compatible topology), or consider the compactification of $\mathcal{X}$. [^1] In that case, the weak topology on $\Delta_{\mathcal{X}}$ is understood with respect to the new topology on $\mathcal{X}$, and may differ from the original one. The real heavy lifting is done by the updating function $f$ being continuous (and assumed to output credal sets). This key property cannot be arbitrarily modified, as shown by the following example.
@@
  • Discontinuity (Hausdorff). Fix any $a\in[0,1-\delta]$ and let $I_{n}=[a,1-\delta-1/n]\to I=[a,1-\delta]$ in $d_{H}$. Then $f(I_{n})=[a+\delta,1-1/n]\to[a+\delta,1]$, while $f(I)=[\max\{0,a-\delta\},1-2\delta]$. Hence $d_{H}\big(f(I_{n}),f(I)\big)\geq\delta$, so $f$ is not Hausdorff-continuous at $\lambda_{2}=1-\delta$.
  • No fixed points. If $\lambda_{2}<1-\delta$, then $f$ shifts the interval right by $\delta$, so equality would force $\delta=0$, a contradiction. If $\lambda_{2}\geq 1-\delta$, then $f([\lambda_{1},\lambda_{2}])=[\max\{0,\lambda_{1}-\delta\},\lambda_{2}-\delta]$, so equality would require $\lambda_{2}=\lambda_{2}-\delta$, again a contradiction. Thus, the set $\mathrm{Fix}(f)$ of fixed points of $f$ is empty.
@@
In applications, one rarely verifies Hausdorff continuity of $f:\mathscr{C}\to\mathscr{C}$ from first principles. Instead, one typically exploits structural properties of the credal sets and of the updating rule, or resorts to numerical diagnostics. We briefly outline three practically relevant situations, followed by an example based on Credal Bayesian Deep Learning, an IPML updating technique based on a generalization of Bayes’ updating.
(1) Finite-dimensional/polyhedral representations. Let $\Delta_{\mathcal{X}}$ be a finite-dimensional simplex, and let credal sets $\mathcal{P}\in\mathscr{C}$ be represented as polytopes, for instance $\mathcal{P}(\eta)=\{p\in\Delta_{\mathcal{X}}:A(\eta)p\leq b(\eta)\}$ or $\mathcal{P}(\eta)=\operatorname{CH}\{p_{1}(\eta),\dots,p_{m}(\eta)\}$, where $\text{CH}(\cdot)$ denotes the convex hull operator, for some parameter $\eta$ encoding vertices or constraints. If the updater is built from operations such as pointwise application of a continuous map $T:\Delta_{\mathcal{X}}\to\Delta_{\mathcal{X}}$, followed by convex hull, $f(\mathcal{P})=\operatorname{CH}\{T(p):p\in\mathcal{P}\}$, or intersection with families of half-spaces whose coefficients depend continuously on $\eta$, then the induced map $\mathcal{P}(\eta)\mapsto f(\mathcal{P}(\eta))$ is continuous with respect to the Hausdorff metric $d_{H}$. Intuitively, small perturbations of the vertices or constraints of $\mathcal{P}$ lead to small perturbations of the vertices or constraints of $f(\mathcal{P})$, and hence to small changes in $f(\mathcal{P})$ in the Hausdorff topology.
(2) Optimization-based updating rules. A second possible pattern is that the updated credal set is defined as the solution set of an optimization problem,
$f(\mathcal{P})=\operatorname*{argmin}_{Q\in\mathcal{A}(\mathcal{P})}\mathfrak{L}(Q),$
where $\mathfrak{L}$ is a loss or divergence functional and $\mathcal{A}(\mathcal{P})$ is a feasible region derived from the initial credal set $\mathcal{P}$ and the observed data. Under standard stability conditions from parametric optimization, for example (i) $\mathfrak{L}$ is continuous (or even strictly convex) in $Q$; (ii) the feasible set $\mathcal{A}(\mathcal{P})$ is nonempty, compact, and depends continuously on $\mathcal{P}$ (e.g., via constraints whose coefficients vary continuously with parameters encoding $\mathcal{P}$); (iii) minimizers are unique or vary continuously as these parameters change, the solution mapping $\mathcal{P}\mapsto f(\mathcal{P})$ is continuous in the Hausdorff metric. Thus, optimization-based updating techniques can inherit Hausdorff continuity from the continuity of their objective and constraints.
(3) Heuristic and numerical diagnostics. Even when an analytic verification is cumbersome, practitioners can still assess the behavior of $f$ heuristically. A simple strategy is:
  1. (1)
    Parametrize credal sets $\mathcal{P}_{\eta}\in\mathscr{C}$ by a finite-dimensional parameter $\eta$ (e.g. bounds, vertices, or hyperparameters) in a normed parameter space.
  2. (2)
    For small perturbations $\kappa\eta$, compute $f(\mathcal{P}_{\eta})$ and $f(\mathcal{P}_{\eta+\kappa\eta})$.
  3. (3)
    Approximate the Hausdorff distance $d_{H}\bigl(f(\mathcal{P}_{\eta}),f(\mathcal{P}_{\eta+\kappa\eta})\bigr)$ by sampling a finite set of distributions from each credal set and computing pairwise distances.
If $d_{H}(f(\mathcal{P}_{\eta}),f(\mathcal{P}_{\eta+\kappa\eta}))$ remains small and scales with $\|\kappa\eta\|$ across a variety of directions in parameter space (and, if relevant, across small perturbations of the data), this provides empirical evidence that the updater behaves in a Hausdorff-continuous way. Conversely, large, irregular jumps in these approximate distances typically signal violations of continuity (for instance caused by hard thresholds, selection of a single “best” model, or conditioning on near-zero likelihood events).
These observations indicate that, in many imprecise probabilistic ML settings of interest, Hausdorff continuity of $f$ is either guaranteed by the way the updater is constructed (polyhedral or optimization-based rules) or can at least be inspected numerically. Let us now turn our attention to Credal Bayesian Deep Learning.
3.1.1. Credal Bayesian Deep Learning (CBDL)